OpenStudy (anonymous):

NEED HELP PRETTY PLEASE;-) Will GIVE MEDALLL AND MESSAGE Rewrite in simplest radical form x^5/6 divided by x^1/6 . Show each step of your process.

iYuko (iyuko):

And a message. lawl

OpenStudy (anonymous):

whatever the fan thing is

OpenStudy (anonymous):

i meant fan I'm new so i don't know what they mean lol

OpenStudy (anonymous):

but really can you help me understand this i remember some of the steps but not all @iYuko

iYuko (iyuko):

How do you start it off?

OpenStudy (anonymous):

you would subtract the two exponents since they have like denominators

iYuko (iyuko):

Okay give me just a second. :)

OpenStudy (anonymous):

so it would be 4/6 and then simplified 2/3 right?

OpenStudy (anonymous):

i think i get it would it be ^ 3 square root and then inside x ^ 2?

iYuko (iyuko):

\[\frac{ x ^{5/6} }{ x ^{1/6} }=x \frac{ 2 }{ 3 }=\sqrt[3]{x ^{2}}\]

OpenStudy (anonymous):

ok i got it right then haha i just needed to think it through logically for a second there.... thanks for the help and re assurance !